OpenStudy (anonymous):

How would I solve for x here? sqrt(3x - 2) + sqrt(11 + x) = 1

Parth (parthkohli):

square both sides.

Parth (parthkohli):

Because you have to get rid of that radical, square both sides.

OpenStudy (anonymous):

right so.. hold on a sec

Parth (parthkohli):

\[\rm 3x - 2 + 2(\sqrt{11+x})(\sqrt{3x - 2}) + 11 + x = 1\]

Parth (parthkohli):

Can you complete it?

OpenStudy (anonymous):

i have 4x + 9 + 2sqrt(3x^2 + 31x - 22) = 1

OpenStudy (anonymous):

then i guess i have to factor

OpenStudy (anonymous):

so I'll just have the sqrt by itself and then put the 4x + 9 on the other side and then I can divide by the 2 in front of the radical to get sqrt(....) = -2x - 4

Parth (parthkohli):

Oh yes, you'd have to factor.

OpenStudy (anonymous):

so I'm on the right track then?

Parth (parthkohli):

I'm trying to find a trick here.

Parth (parthkohli):

No solutions!

OpenStudy (anonymous):

that's correct

Parth (parthkohli):

I solved it and got no solutions as the answer.

OpenStudy (anonymous):

how?

Parth (parthkohli):

I actually raised both sides to power of 3. ;)

Parth (parthkohli):

But I did it on paper and that was fast...

OpenStudy (anonymous):

why did you raise both sides to the third?

Parth (parthkohli):

I was just trying and testing.

Parth (parthkohli):

... and then that didn't work.

Parth (parthkohli):

so I changed my route again.

Parth (parthkohli):

Let me show my solution here:\[\rm \sqrt{3 x - 2} + \sqrt{x + 11} = 1\]\[\rm 3x - 2 + 2\cdot\sqrt{3x - 2}\cdot \sqrt{x + 11} + x+11 = 1 \]\[\rm 4x + 9 + 2\cdot \sqrt{3x - 2}\cdot \sqrt{x + 11} = 1\]\[\rm 4x + 9 + 2\sqrt{3x^2 + 31x - 22} = 1\]\[\rm 4x + 2\sqrt{3x^2 + 31x - 22} = -8\]\[\rm 2x + \sqrt{3x^2 + 31x - 22}=-8\]

Parth (parthkohli):

Square both sides again.

OpenStudy (anonymous):

Dont need to, isolate the square root an then do it, the square root vanishes

Parth (parthkohli):

Oh, good point

Parth (parthkohli):

\[\rm \sqrt{3x^2 + 31x - 22}= -2x - 8\]\[\rm 3x^2 + 31x - 22 = 4x^2 +32x+64\]

Parth (parthkohli):

I mistakenly squared both sides on paper but it still worked.